OpenStudy (lisa123):

Find all solutions to the equation 2sin theta- srt(3)=0 Write your answer in radians in terms of π, and use the "or" button as necessary

OpenStudy (anonymous):

sintheta=sqrt(3)/2

OpenStudy (anonymous):

so theta=sin-1(sqrt(3)/2)

OpenStudy (anonymous):

theta=60 degrees.

OpenStudy (anonymous):

to convert 60 deg to radians, you multiply by pi/180

OpenStudy (anonymous):

no need to convert, ignore degrees entirely

OpenStudy (anonymous):

\[\sin(\theta)=\frac{\sqrt3}{2}\to \theta=\frac{\pi}{3}\] no degrees where endangered in this process

OpenStudy (anonymous):

lol its good for him to learn how to convert tho..

OpenStudy (anonymous):

course that aint all of them that i just one there are an infinite number of solutions

OpenStudy (lisa123):

so the solutions are sqrt(3)/2 and pi/3?

OpenStudy (anonymous):

there is another on the other side of the unit circle, across the y axis that would be \(\frac{2\pi}{3}\) and then there are more

OpenStudy (anonymous):

oh no dear lets to slow

OpenStudy (anonymous):

yesh so it should be (pi/3)+2pi*k or (2pi/3))+2pi*k

OpenStudy (anonymous):

where k is any integer constant

OpenStudy (lisa123):

alright thank you guys so much!